OpenStudy (anonymous):

The price of a pair of skis at the Solitude Ski Shop is $254. This price includes the store's cost for the skis plus a markup at the rate of 27%. Find Solitude's cost for the skis.

OpenStudy (lilg132):

322.58

OpenStudy (farmdawgnation):

So... the problem you're describing is a solve-for-x situation, where x is the cost of the skis. Setting up this equation should get you the answer. \[254 = x + .27(x)\]

OpenStudy (lilg132):

254 / 100 = 2.54 (this is 1 percent of the price) 2.54 * 27 = 68.58 (this is 27 percent of the price) 68.58 + 254 = 322.58 (this is the final price)

OpenStudy (anonymous):

that was wrong

OpenStudy (farmdawgnation):

lilg132 -- No, the price they are selling for is $254. It doesn't make sense that the store would buy the skis at $322.58 and then sell them for $254.

OpenStudy (anonymous):

right... it still was counted wrong

OpenStudy (anonymous):

i got one more chance lol

OpenStudy (farmdawgnation):

Ok, so I didn't set this up correctly the first time. Look at it like this.... $254 = 127% of the sale price. x = 100% of the sale price. Therefore: \[(254\div127)=(x \div 100)\] Solve for x. :)